Problem 2-27 (Algorithmic) In an article about investment alternatives, Money magazine reported that drug stocks provide a potential for long-term growth, with over 50% of the adult population of the United States taking prescription drugs on a regular basis. For adults age 65 and older, 86% take prescription drugs regularly. For adults age 18 to 64, 48% take prescription drugs regularly. The 18-64 age group accounts for 81.7% of the adult population (Statistical Abstract of the United States, 2008). a. What is the probability that a randomly selected adult is 65 or older? If required, round your answer to three decimal places. P (A) b. Given that an adult takes prescription drugs regularly, what is the probability that the adult is age 65 or older? If required, round your answer to four decimal places. P (AlD)-
Answer #1

a)
probability that a randomly select adult is 65 and older =1 - 0.817 = 0.183

b)
A - adult takes prescription drugs
D - 65 or older

P(A |D) = P(A and D)/P(D)
=   (0.183 * 0.86 )/(0.183 * 0.86 + 0.817 * 0.48 )
= 0.2863849

=0.2864